OpenStudy (anonymous):

Which of the following is a polynomial with roots sqrt -3 , sqrt 3 , and -2? a) x3 - 2x2 - 3x + 6 b) x3 + 2x2 - 3x - 6 c) x3 - 3x2 - 5x + 15 d) x3 + 3x2 - 5x - 15

OpenStudy (anonymous):

@Ahsome do you know this? Could you help me with this one

OpenStudy (anonymous):

@ganeshie8 could you please help me?

OpenStudy (ahsome):

You could try dividing all those equations by their roots and see if it fits. Example with the first one: \[\frac{x^3 - 2x^2 - 3x + 6}{\sqrt{-3}}\] \[\frac{\text{Above Answer}}{\sqrt{3}}\] \[\frac{\text{Above Answer}}{-2}\] And see if you can find an answer? If at any point it doesn't work, simply change the top equation\((x^3 - 2x^2 - 3x + 6)\) with the other possibility

OpenStudy (anonymous):

try the roos given..... if the answer is zero then this is the eq.

OpenStudy (anonymous):

Thank you! @Ahsome @AJ01

OpenStudy (ahsome):

No problem @natgirl200 :D

OpenStudy (ahsome):

Psst, easy trick. Simply use your scientific calculator to find the roots of the equation and see if they match up, Then, if they do, do it manually and say you are so lucky you keep getting it the first time ;)

OpenStudy (anonymous):

Oh okay thanks!! I got the answer B :) @Ahsome

OpenStudy (ahsome):

Great job :)